# arXiv:2603.01473v1[astro-ph.GA] 02 Mar 2026

### 02 Mar 2026 v1

@@ Abstract
Little red dots (LRDs) draw extensive attention because of their unique observational characteristics and apparent overabundance in the early Universe, raising new insights into early black hole formation and growth. Early studies show that LRDs exhibit weak variability in broad-band photometry and emission-line fluxes, suggesting a preference for super-Eddington accretion or disfavouring an AGN origin. However, the cadence of the current data, and therefore, the resulting light curves for LRDs, is limited, preventing us from placing strong constraints on their variability. Based on Zwicky Transient Facility (ZTF) light curves with a baseline of $\sim 6$ years, we here study the optical variability of seven previously reported local analogs of LRDs at $z\sim 0.3$, offering an insight into LRDs from a low-redshift sample. Three out of seven local analogs show excess variances on all three bands of their light curves, and two of them can be fitted with the damping random walk model, supporting their AGN origins for the variability. The remaining sources show weak variance in at least one band, but no detectable variability at the current sensitivity level, exhibiting $\rm SF_{\infty}$ upper limits consistent with estimates from high-redshift (high-$z$) LRDs. Their non-detection of variability is likely due to the large photometric uncertainty. As an implication, by simulating long baseline light curves with the variability amplitude of local analogs and adopting JWST observation cadence, we investigate the limitation of the variability amplitude estimate for LRDs. Our mock observations imply that the current constraints on LRDs’ variability are probably underestimated. This underestimation might be induced by the short temporal baseline of observations, as well as the intrinsic scatter of the empirical $M_{\rm BH}-\tau$ relation.
@@
facilities: ZTFsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), celerite (Foreman-Mackey et al., 2017; Foreman-Mackey, 2018), emcee (Foreman-Mackey et al., 2013), grizli (Brammer, 2023a), msaexp (Brammer, 2023b),
@@ II.1 Low-mass AGNs and Local LRDs in GPs
Subsequently, Lin et al. (2025a) identified seven local analogs of LRD within this sample, based on similar observational characteristics, i.e., V-shaped UV-optical SED, broad-line emission, and compact morphology. In Figure 2 we show an example of their SEDs covering rest-frame wavelength from $\sim 0.1$ to $20\,\mu m$. For comparison, we also show the best-match LRD’s SED (PRIMER; PI: Dunlop and RUBIES; PI: de Graaff.) in red, which was reported in Setton et al. (2025).
@@ IV AGN Origin of Local LRDs
Figure 6: Examples of mock observations with temporal offsets of 100 (blue), 620 (brown), 1300 (green), and 1900 (purple) days. Top: The black curve shows a DRW light curve with $\tau_{\rm DRW}$ of 144 days and $\rm SF_{\infty}$ of 0.18 mag. For each offset, the rest-frame sampling epochs (0, 22, 63, and 87 days) are adopted from those of the LRD UDS-47509 (Kokubo and Harikane, 2025). Bottom: The posterior distributions of $SF_{\infty}$. The black lines mark the 90th-percentile upper limits.
@@ V.1 Mock Observation
We first construct 50 realizations of high-cadence model light curves spanning 2500 days based on the DRW model. We adopt the solution given in Equation (10) of Moreno et al. (2019), assuming an $\rm SF_{\infty}$ of 0.18 mag (equal to that of J092834+292136) and a intrinsic timescale $\tau_{\rm int}$ of 144 days, derived from the $M_{\rm BH}-\tau$ relation using the black hole mass of LRD UDS-47509.
For each realization, we evaluate the model fluxes at the rest-frame epochs (0, 22, 62, and 87 days), corresponding to the cadence of the JWST epoch pairs for UDS-47509 (Kokubo and Harikane, 2025). This procedure generates a set of sampled light curves that mimic the JWST LRD observations. To ensure consistency with high-$z$ studies, we derive the $\rm SF_{\infty}$ posterior and estimate the upper limit of $\rm SF_{\infty}$ assuming a decorrelation timescale based on the empirical $M_{\rm BH}-\tau$ relation from Wang et al. (2023) for dwarf AGNs. Specifically, we adopt 144 days for UDS-47509 with $M_{\rm BH}=10^{8.34}\,M_{\odot}$, which matches the intrinsic timescale of simulated light curves. Four illustrative examples are shown in Figure 6.

### 16 Mar 2026 v2

# Local Analogs of Little Red Dots: Optical Variability and Evidence for an AGN Origin

@@ Abstract
Little red dots (LRDs) draw extensive attention because of their unique observational characteristics and apparent overabundance in the early Universe, raising new insights into early black hole formation and growth. Early studies show that LRDs exhibit weak variability in photometry and emission-line fluxes, suggesting a preference for super-Eddington accretion or disfavouring an AGN origin. However, the current data is limited, preventing us from placing strong constraints on their variability. Based on ZTF light curves with a baseline of $\sim 6$ years, we here study the optical variability of seven previously reported local analogs of LRDs at $z\sim 0.3$, offering an insight into LRDs from a low-redshift sample. Three out of seven local analogs show excess variances on all three bands of their light curves, and two of them can be fitted with the damping random walk model, supporting their AGN origins for the variability. The remaining sources show weak variance in at least one band, but no detectable variability, exhibiting $\rm SF_{\infty}$ upper limits consistent with estimates from high-redshift LRDs. Their non-detection of variability is likely due to the large photometric uncertainty. As an implication, by simulating long baseline light curves with the variability amplitude of local analogs and adopting JWST observation cadence, we investigate the limitation of the variability amplitude estimate for LRDs. Our mock observations imply that the current constraints on LRDs’ variability are probably underestimated. This underestimation might be induced by the short temporal baseline of observations, as well as the intrinsic scatter of the empirical $M_{\rm BH}-\tau$ relation.
@@
facilities: ZTF (Masci et al., 2019)software: ZTF Lightcurves (ZTF Team, 2025), astropy (Astropy Collaboration et al., 2013, 2018, 2022), celerite (Foreman-Mackey et al., 2017; Foreman-Mackey, 2018), emcee (Foreman-Mackey et al., 2013), grizli (Brammer, 2023a), msaexp (Brammer, 2023b),
@@ II.1 Low-mass AGNs and Local LRDs in GPs
Subsequently, Lin et al. (2025a) identified seven local analogs of LRD within this sample, based on similar observational characteristics, i.e., V-shaped UV-optical SED, broad-line emission, and compact morphology. In Figure 2 we show an example of their SEDs covering rest-frame wavelength from $\sim 0.1$ to $20\,\mu m$. For comparison, we also show the best-match LRD’s SED (JADES; PID 1181, PI: Eisenstein) in red, which was reported in Setton et al. (2025).
@@ IV AGN Origin of Local LRDs
Figure 6: Examples of mock observations with temporal offsets of 100 (blue), 620 (brown), 1300 (green), and 1900 (purple) days. Top: The black curve shows a DRW light curve with $\tau_{\rm DRW}$ of 144 days and $\rm SF_{\infty}$ of 0.18 mag. For each offset, the rest-frame sampling epochs (0, 22, 63, and 87 days) are adopted from those of the LRD MSAID38108 (Kokubo and Harikane, 2025). Bottom: The posterior distributions of $SF_{\infty}$. The black lines mark the 90th-percentile upper limits.
@@ V.1 Mock Observation
We first construct 50 realizations of high-cadence model light curves spanning 2500 days based on the DRW model. We adopt the solution given in Equation (10) of Moreno et al. (2019), assuming an $\rm SF_{\infty}$ of 0.18 mag (equal to that of J092834+292136) and a intrinsic timescale $\tau_{\rm int}$ of 144 days, derived from the $M_{\rm BH}-\tau$ relation using the black hole mass of LRD MSAID38108.
For each realization, we evaluate the model fluxes at the rest-frame epochs (0, 22, 62, and 87 days), corresponding to the cadence of the JWST epoch pairs for MSAID38108 (Kokubo and Harikane, 2025). This procedure generates a set of sampled light curves that mimic the JWST LRD observations. To ensure consistency with high-$z$ studies, we derive the $\rm SF_{\infty}$ posterior and estimate the upper limit of $\rm SF_{\infty}$ assuming a decorrelation timescale based on the empirical $M_{\rm BH}-\tau$ relation from Wang et al. (2023) for dwarf AGNs. Specifically, we adopt 144 days for MSAID38108 with $M_{\rm BH}=10^{8.34}\,M_{\odot}$, which matches the intrinsic timescale of simulated light curves. Four illustrative examples are shown in Figure 6.